OpenStudy (anonymous):

someone plz help me solve complete the ordered triple 4a+3b+c=-9 a-3b+2c=27 11a-2b+3c=28 step by step so i can learn how to do on my own i keep getting the wrong answer

OpenStudy (anonymous):

subtract twice equation 1 from equation 2 and thrice equation 1 from equation 3. that should eliminate c and you will have two new equations with just a and b. solve for a and b and substitute their values in one of the original equations to get c

OpenStudy (anonymous):

your suppose to add them not subtract i thought

OpenStudy (anonymous):

no you can manipulate them any way you like. subtraction is simply negative adding. 4-2 is 4+(-2). Think about it this way then add negative twiceequation 1 from equation 2 and negative thrice equation 1 from equation 3. that should eliminate c and you will have two new equations with just a and b. solve for a and b and substitute their values in one of the original equations to get c

OpenStudy (anonymous):

http://www.flickr.com/photos/62007598 @N03/5645382314/

OpenStudy (anonymous):

I'm not good with tye , writing more easy for me

OpenStudy (anonymous):

it says its a bad link

OpenStudy (anonymous):

you can see?

OpenStudy (anonymous):

I see very good

OpenStudy (anonymous):

it says page not found oops looks like you followed a bad link

OpenStudy (anonymous):

let tye

OpenStudy (anonymous):

...........4a=3b+c=-9 ...........a-3b+2c=27 Add....5a = 3c =18 1st + 3rd time 2 to the frist equation 8a+6b+2c=-18 time (3)*3rd rquation 33a-6b+9c=84 ...............8a+6b+2c=-18 ..............33a-6b+9c=84 add...........a+11c=66 time(11) to 5a+3c=18 ..............55a+33c=198 time(-3) to 41a +11c=66 ...............-123a-33c=-198 ................55a+33c=198 ...............-123a-33c=-198 add...........-69a=0 ......................a=0 sutitution a=0 find 5(0)=3c=18 ...........c=6 sutitution c=6 find find b 11(0)-2b+3(6)=28 0 -2b+18=28 ...........-2b=28-18 ...............b=10/-2 ...............b=-5 a=0...., ....5,......c= 6

OpenStudy (anonymous):

Add....5a = 3c =18 I tye wrong: 5a+3c=18 I told you I'm not good tye you check I may tye wrong

OpenStudy (anonymous):

a=0 b=-5 c=6

OpenStudy (anonymous):

...........4a=3b+c=-9 ...........a-3b+2c=27 Add....5a +3c =18 1st + 3rd time 2 to the frist equation 8a+6b+2c=-18 time (3)*3rd rquation 33a-6b+9c=84 ...............8a+6b+2c=-18 ..............33a-6b+9c=84 add........41a+11c=66 time(11) to 5a+3c=18 ..............55a+33c=198 time(-3) to 41a +11c=66 ...............-123a-33c=-198 ................55a+33c=198 ...............-123a-33c=-198 add...........-69a=0 ......................a=0 sutitution a=0 find 5(0)=3c=18 ...........c=6 sutitution c=6 find find b 11(0)-2b+3(6)=28 0 -2b+18=28 ...........-2b=28-18 ...............b=10/-2 ...............b=-5 a=0...., ..b=-,......c= 6 I check every thing is correct now I have to go now bye

OpenStudy (anonymous):

you still here

OpenStudy (anonymous):

check again http://www.flickr.com/photos/62007598 @N03/5645382314/

OpenStudy (anonymous):

http://img857.imageshack.us/i/mathk.jpg/